# MI-HET-855

### v1

@@ Abstract
No ($m_{\phi},m_{Z^{\prime}}$) [MeV] $g$ [GeV-1] $\nu$-MB $\bar{\nu}$-MB $\mu$B
1 (45.41, 48.49) $3.93\times 10^{-7}$ 2.02 1.49 1.13
2 (87.67, 93.63) $8.20\times 10^{-7}$ 1.76 1.56 1.20
3 (119.08, 114.06) $1.06\times 10^{-6}$ 1.76 1.58 1.28
4 (193.07, 138.95) $1.55\times 10^{-6}$ 1.76 1.66 1.35
Table 1: Summary of MiniBooNE and MicroBooNE fits for various $m_{\phi}$ and $m_{Z^{\prime}}$ values ($m_{Z^{\prime}}=3m_{\chi}$). The effective coupling $g$ is given by $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{q}$.
Results. A crucial factor determining the photon signal at MiniBooNE is the relative mass gap between $\phi$ and $Z^{\prime}$. Although the MiniBooNE excess peaks at a lower energy ($E_{\gamma}\sim 200-300$ MeV), it exhibits a long tail extending up to 1.2 GeV. Similarly, while the MiniBooNE cosine excess peaks at $\cos\theta=1$, a prominent off-forward component is also observed. Given that neutrinos or dark matter at MiniBooNE peak around 500 MeV, $m_{Z^{\prime}}$ must be sufficiently large to enable significant energy transfer to the photon—yet not so large that it transfers nearly all of the energy. Additionally, the scalar mediator exchanged between the $Z^{\prime}$-$\gamma$ system and the nucleus must not be much heavier than the $Z^{\prime}$; otherwise, the energy would predominantly be transferred to the nucleons or nucleus. At the same time, the scalar must be heavy enough to allow for momentum transfer to the nucleus, facilitating the observed off-forward photon emission.
Figure 2 displays an example MiniBooNE and MicroBooNE fit ( 2 in Table 1), incorporating contributions from both dark matter and neutrinos, as $Z^{\prime}$ couples to both $\chi$ and $\nu$. Furthermore, we find that the same parameter choices yield the same $\chi^{2}$/dof when fitting the excess with neutrino and dark matter signals separately (see Fig. 4 of Appendix E). We also identify other combinations of $m_{Z^{\prime}}$, $m_{\phi}$, and $g$ that provide fits to the energy and cosine excesses observed at MiniBooNE, yielding similar values of $\chi^{2}$/dof. These parameters are shown in Table 1 along with the average $\chi^{2}$/dof (averaged over the cosine and energy fits). We find that mediators with comparable masses in the range of $50\leavevmode\nobreak\ \text{MeV}\lesssim m_{\phi/Z^{\prime}}\lesssim 200\leavevmode\nobreak\ \text{MeV}$ provide good fits to the MiniBooNE excess, with an average $\chi^{2}$/dof between 1 and 2. Additionally, we observe that the dark matter flux and the $2\to 3$ cross section change only minimally with variations in dark matter mass, as long as $2m_{\chi}<m_{Z^{\prime}}$. This minimal sensitivity of the $\chi^{2}$/dof values to $m_{\chi}$ is further demonstrated in Table 2 of Appendix E.
Although the fits were determined by minimizing $\chi^{2}$ only over MiniBooNE’s neutrino-mode data, we find that these same parameters yield $\chi^{2}$/dof values at MicroBooNE that are lower than $\chi^{2}$/dof obtained under the null hypothesis—i.e., assuming SM events only. We also find that the MicroBooNE fits do not reproduce the forward and lower energy excess as effectively as the fits for MiniBooNE. This is mainly due to the choice of flavor-dependent Yukawa couplings, which results in reduced coherent contributions at the MicroBooNE detector. However, the $\chi^{2}$/dof values in MicroBooNE in Table 1 suggest that smaller $m_{\phi}$ and $m_{Z^{\prime}}$ values are preferred to reasonably accommodate both MiniBooNE and MicroBooNE excesses under the current Yukawa coupling choices. Given the significantly lower statistics in MicroBooNE compared to MiniBooNE’s neutrino mode, we refrain from drawing quantitative conclusions based on the MicroBooNE fits.
Experimental Constraints. Existing searches from experiments such as NA62, PIENU, BaBar, CHARM, and E787 would constrain the mass and coupling of the mediators in our model. We first identify the dominant decay modes of the mediators $\phi$ and $Z^{\prime}$ to determine the relevance of each constraint.
Since the MiniBooNE fits prefer a gauge boson that is slightly heavier than the scalar, the $Z^{\prime}$ is kinematically allowed to decay to $\bar{\nu}\nu$, $\bar{\chi}\chi$, and $\phi\gamma$. Since $g_{\chi}=2.5$ is much larger than $g_{\nu}$ and $g_{\phi Z^{\prime}\gamma}$, $Z^{\prime}$ decays predominantly into dark matter with a branching ratio of $\sim 100\%$. For scalars of $\mathcal{O}(100\leavevmode\nobreak\ \text{MeV})$ that are lighter than $Z^{\prime}$, the allowed decay modes include $\phi\to\gamma\chi\bar{\chi}$, $\phi\to\gamma\nu\bar{\nu}$, and $\phi\to\gamma\gamma$; the first two processes are facilitated by an off-shell $Z^{\prime}$ and the third is through quark loops. The branching ratio for $\phi\to\gamma\gamma$ cannot be determined solely by the $u$ and $d$ charges, as the presence of other Yukawa couplings also affects the decay width. Since we do not delve into the intricacies of the scalar model, we neglect the bounds that involve the decay mode $\phi\rightarrow\gamma\gamma$. Due to $g_{\chi}\gg g_{\nu}$, the dominant decay mode of the scalar is $\phi\to\gamma\chi\bar{\chi}$. We also note that when $m_{Z^{\prime}}<m_{\phi}$, the scalar dominantly decays into $\phi\to\gamma Z^{\prime}$ instead.
While $Z^{\prime}$ can be produced via its coupling to neutrinos, they leave only invisible signatures. Therefore, the strongest constraints on $g_{\nu}$ are from the upper limit on the charged-meson decays as measured by PIENU PIENU:2021clt ($\pi^{\pm}\rightarrow e^{\pm}\nu_{e}Z^{\prime}/\mu^{\pm}\nu_{\mu}Z^{\prime}$) and NA62 NA62:2021bji ($K^{\pm}\rightarrow\mu^{\pm}\nu_{\mu}Z^{\prime}$). These bounds are summarized in Fig. 5 in Appendix. F where the upper bound on $g_{\nu}$ for $10\leavevmode\nobreak\ \text{MeV}\leq m_{Z^{\prime}}\leq 200\leavevmode\nobreak\ \text{MeV}$ is $2\times 10^{-2}\leq g_{\nu}\leq 6\times 10^{-2}$.
The dimension-5 couplings $g_{\phi Z^{\prime}\gamma}$ can arise as a result of UV models that contain $Z^{\prime}$ and $\phi$ couplings to charm and bottom quarks. For gauge and Yukawa couplings where $g_{Z^{\prime}c/b}\times y_{\phi c/b}\sim\pm\mathcal{O}(0.1)$, the integration over quark loops can give rise to dimension-5 couplings that are $\mathcal{O}(10^{-3})$. The leading bounds on the dimension-5 coupling, $g_{\phi Z^{\prime}\gamma}>10^{-4}\leavevmode\nobreak\ \text{GeV}^{-1}$, are from monophoton searches at BaBar BaBar:2001yhh ; BaBar:2013agn and CHARM Gninenko:2011uv ; Gninenko:2012eq , and invisible searches at E949/E787 E787:2002qfb ; E787:2004ovg ; E949:2007xyy ; BNL-E949:2009dza . A single-photon final state arises in the $e^{+}e^{-}$ collider via the processes $e^{+}e^{-}\rightarrow\gamma^{*}\rightarrow Z^{\prime}\phi$ where the scalar decays via $\phi\to\gamma\chi\bar{\chi}$ or $\phi\to\gamma Z^{\prime}$ deNiverville:2018hrc within $\sim 1$ m. CHARM, a proton-on-target experiment, can probe the model via $\pi^{0}/\eta\rightarrow\gamma(\gamma^{*}\rightarrow Z^{\prime}\phi)$ Gninenko:2011uv ; Gninenko:2012eq ; CHARM:1983ayi followed by $\phi\to\gamma\bar{\chi}\chi$ at the detector 480 m away. Complementary to the above searches, E949/E787 looks for scalars produced by $K^{+}\rightarrow\pi^{+}X$ when $X=\text{invisible}$ E787:2002qfb ; E787:2004ovg ; E949:2007xyy ; BNL-E949:2009dza where “invisible” constitutes the scalar decaying into $\gamma Z^{\prime}/({Z^{\prime}}^{*}\rightarrow\bar{\chi}\chi)$ outside the E949/E787 detector which is $\sim 1.45\leavevmode\nobreak\ \text{m}$ in size. Figure 3 shows these bounds for various mass gaps between the initial and final states, which is defined as $\Delta m=m_{\phi}-2m_{\chi}$ for $2m_{\chi}<m_{\phi}<m_{Z^{\prime}}$, and $\Delta m=m_{\phi}-m_{Z^{\prime}}$ for $m_{\phi}>m_{Z^{\prime}}$. Since the decay width is smaller for smaller mass gaps, the lifetimes of $\phi$ are generally longer. Therefore, the typical couplings that can be excluded for smaller mass gaps are larger than in the case of larger mass gaps. We also see that the energy of the photons produced at BaBar are less than BaBar’s energy thresholds, 2.2 GeV, for $\Delta m/m_{\phi}\lesssim 0.4$. Hence, the BaBar constraints are relevant for larger mass gaps, as depicted in the $\Delta m/m_{\phi}=0.6$ scenario.
The ongoing Belle-II Belle-II:2010dht experiment can probe couplings lower than BaBar due to its larger luminosity and lower photon thresholds. Due to excessive monophoton backgrounds and the lack of monophoton triggers BelleIIMW , there are no existing monophoton limits at Belle-II. However, we make preliminary, ballpark predictions for Belle-II by assuming that the monophoton backgrounds are comparable to those for three-photon events belle2 . The green dashed lines in Fig. 3 depict our rather optimistic predictions for a total luminosity of $50\leavevmode\nobreak\ \text{ab}^{-1}$ with $N_{\text{bkg}}=7\times 10^{5}$. [^3] In reality, we expect monophoton backgrounds to be much larger than that of three photons. Therefore, the limits presented would scale up by $\sqrt{N_{1\gamma}^{\text{true}}/(7\times 10^{5})}$, thus relaxing the upper bound.
Quark-philic scalars are constrained by the two-body branching ratio of $K^{+}\rightarrow\pi^{+}\pi^{0}$ ParticleDataGroup:2024cfk , where values satisfying $\text{BR}(K^{+}\rightarrow\pi^{+}\phi)<0.2067$ are allowed. Additional constraints arise from MAMI A2atMAMI:2014zdf ; CrystalBallatMAMI:2010slt ; Oset:2008hp , which limit the branching ratio of eta mesons via $\eta\rightarrow\pi^{0}\gamma\gamma<3\times 10^{-4}$, imposing the condition $y_{q}^{2}\sum_{q}N_{C}Q_{q}^{2}x_{q}/m_{q}\leq 0.2g_{\phi Z^{\prime}\gamma}$. However, these bounds depend on the Yukawa couplings to strange and other heavy quarks. Therefore, Yukawa couplings of $\mathcal{O}(0.1)$ remain viable.
The recent analysis from MicroBooNE constrains the neutrino flux-averaged coherent cross section to obey $\sigma<1.49\times 10^{-41}\leavevmode\nobreak\ \text{cm}^{2}$ MicroBooNE:2025rsd . However, the distribution of single photons as a function of true $E_{\gamma}$ and $\cos\theta_{\gamma}$ for the $2\to 3$ scattering differs from that used in Ref. MicroBooNE:2025rsd . In particular, the $2\to 3$ BSM process we consider here gives rise to photons that are more energetic than those from the SM $\text{NC}1\gamma$ process, which arise from $Z$-boson exchange with the nucleus that induces a ground-state-to-ground-state transition in the nucleus with an associated photon PhysRevC.89.015503 . Therefore, this cross section constraint only weakly applies, with a $\sim$10overlap in the true photon energy and angular distribution template in Ref. MicroBooNE:2025rsd to the coherent part of the $2\to 3$ process. We find that 10% of the coherent $2\to 3$ cross section satisfies the upper bound of $1.49\times 10^{-41}\leavevmode\nobreak\ \text{cm}^{2}$.
Based on these constraints, we find that there exist coupling values capable of fitting the excess with all best-fit points listed in Table 1. These are projected on the $(m_{\phi},g_{\phi Z^{\prime}\gamma})$ parameter space in Fig. 3 including $m_{\chi}$ choices other than $m_{Z^{\prime}}=3m_{\chi}$. The example in Fig. 2 corresponds to 2 in the middle panel of Fig. 3, and the values $g_{\nu}=3\times 10^{-3}$, $g_{\phi Z^{\prime}\gamma}=10^{-3}\leavevmode\nobreak\ \text{GeV}^{-1}$, and $y_{q}=0.27$ provide a consistent fit.
Figure 3: Bounds on $g_{\phi Z^{\prime}\gamma}$ as a function of $m_{\phi}$ from BaBar and CHARM monophoton searches, and E787/E949 invisible searches for different $\Delta m$ (see text for more details). The green dashed line shows the predictions for Belle-II at 50 ab-1. We also mark the four best-fit points from Table 1 in blue circles with several $m_{\chi}$ choices.
@@
We performed fits to the energy and angular distributions of the excess events observed by MiniBooNE and MicroBooNE and found that the required mediator and dark matter masses and couplings are consistent with current experimental bounds. We considered $Z^{\prime}$ couples with both neutrino and dark matter in our example to produce it using the neutrino leg of the 3-body charged-pion decays, which need not be true to obtain the dark matter fit. For mediator masses in the range $50\leavevmode\nobreak\ \text{MeV}\lesssim m_{\phi/Z^{\prime}}\lesssim 200\leavevmode\nobreak\ \text{MeV}$ in our scenario, we find good fits to the MiniBooNE data, with the average $\chi^{2}$/dof values lying between 1 and 2. These scenarios can be further probed in ongoing experiments such as ICARUS, SBND, and Belle-II. Importantly, the $2\to 3$ scattering mechanism plays a crucial role in explaining the excess: vector exchange with the nucleus tends to produce forward-peaked photons, while pseudoscalar exchange leads to off-forward photon emission due to the Lorentz structure of the interaction with the nucleus.
@@
Acknowledgements. We would like to thank Kevin J. Kelly and Vishvas Pandey for their helpful discussions. The work of BD and AK is supported by the U.S. DOE Grant DE-SC0010813. AT acknowledges support in part by the DOE grant DE-SC0010143.

### v2

@@ Abstract
1 (45.41, 48.49) $5.9\times 10^{-7}$ 2.02 1.49 1.13
2 (87.67, 93.63) $9.0\times 10^{-7}$ 1.76 1.56 1.20
3 (119.08, 114.06) $1.11\times 10^{-6}$ 1.75 1.58 1.28
4 (193.07, 138.95) $1.55\times 10^{-6}$ 1.75 1.66 1.35
Table 1: Summary of the $\chi^{2}$ values of the MiniBooNE fits and MicroBooNE overlays for various $m_{\phi}$ and $m_{Z^{\prime}}$ values ($m_{Z^{\prime}}=3m_{\chi}$). The effective coupling for dark matter fits are given in terms of the couplings required for scattering, i.e., $g_{\phi Z^{\prime}\gamma}y_{\phi}$, and the neutrino fits are given by $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{\phi}$.
Results. A crucial factor determining the photon signal at MiniBooNE is the relative mass gap between $\phi$ and $Z^{\prime}$. Although the MiniBooNE excess peaks at a lower energy ($E_{\gamma}\sim 200-300$ MeV), it exhibits a long tail extending up to 1.2 GeV. Similarly, while the MiniBooNE cosine excess peaks at $\cos\theta=1$, a prominent off-forward component is also observed. Given that neutrinos or dark matter at MiniBooNE peak around 500 MeV, $m_{Z^{\prime}}$ must be sufficiently large to enable significant energy transfer to the photon, yet not so large that it transfers nearly all of the energy. Additionally, the scalar mediator exchanged between the $Z^{\prime}$-$\gamma$ system and the nucleus must not be much heavier than the $Z^{\prime}$; otherwise, the energy would predominantly be transferred to the nucleons or nucleus. At the same time, the scalar must be heavy enough to allow for momentum transfer to the nucleus, facilitating the observed off-forward photon emission.
Figure 2 displays an example MiniBooNE fit along with the MicroBooNE overlay ( 2 in Table 1), for dark matter scattering. We also identify other combinations of $m_{Z^{\prime}}$, $m_{\phi}$, and $g$ that provide fits to the energy and cosine excesses observed at MiniBooNE, yielding similar values of $\chi^{2}$/dof. These parameters are shown in Table 1 along with the average $\chi^{2}$/dof (averaged over the cosine and energy fits). We find that mediators with comparable masses in the range of $50\leavevmode\nobreak\ \text{MeV}\lesssim m_{\phi/Z^{\prime}}\lesssim 200\leavevmode\nobreak\ \text{MeV}$ provide good fits to the MiniBooNE excess, with an average $\chi^{2}$/dof between 1 and 2. Table 1 also contains the fit parameters for neutrino-only scattering. While the dark matter flux’s mean energy is less than that of the neutrino flux, the energy flux above 600 MeV, which is the main contributor to the $2\to 3$ cross section, has a similar trend. Therefore, within the errors of the measurements, the shape of the neutrino excess is similar to that of dark matter shown in Fig. 2, consequently leading to qualitatively similar $\chi^{2}$ values. Furthermore, we find that the same parameter choices yield the same $\chi^{2}$/dof when fitting the excess by combining neutrino and dark matter signals (see Appendix. E), assuming that dark matter is produced via the same neutrinophilic $Z^{\prime}$ that facilitates neutrino $2\to 3$ scattering. We further observe that the dark matter flux and the $2\to 3$ cross section change only minimally with variations in dark matter mass, as long as $2m_{\chi}<m_{Z^{\prime}}$. This minimal sensitivity of the $\chi^{2}$/dof values to $m_{\chi}$ is demonstrated in Appendix. E.
Although the fits were determined by minimizing $\chi^{2}$ only over MiniBooNE’s neutrino-mode data, we find that these same parameters yield $\chi^{2}$/dof values at MicroBooNE that are lower than $\chi^{2}$/dof obtained under the null hypothesis, i.e., SM events only hypothesis. For reference, the average $\chi^{2}/\text{dof}$ for the MiniBooNE neutrino mode, anti-neutrino mode, and MicroBooNE under the SM-only hypothesis are 4.83, 2.02, and 1.69, respectively [^2] The MiniBooNE values are for our assumed systematics. One must include the covariance matrix to match the official results. We also find that the new physics signal at MicroBooNE does not reproduce the forward and lower energy excess as effectively as the fits for MiniBooNE. This is mainly due to the choice of flavor-dependent Yukawa couplings, which result in reduced coherent contributions at the MicroBooNE detector. However, the $\chi^{2}$/dof values in MicroBooNE in Table 1 suggest that smaller $m_{\phi}$ and $m_{Z^{\prime}}$ values, such as fit 1, are preferred to reasonably accommodate both MiniBooNE and MicroBooNE excesses under the current Yukawa coupling choices. Given the significantly lower statistics in MicroBooNE compared to MiniBooNE’s neutrino mode, we refrain from drawing quantitative conclusions based on the MicroBooNE overlays.
Constraints. The recent analysis from MicroBooNE constrains the neutrino flux-averaged coherent cross section to obey $\sigma<1.49\times 10^{-41}\leavevmode\nobreak\ \text{cm}^{2}$ MicroBooNE:2025rsd . However, the distribution of single photons as a function of true $E_{\gamma}$ and $\cos\theta_{\gamma}$ for the $2\to 3$ scattering differs from that used in Ref. MicroBooNE:2025rsd . As shown in Appendix F, the $2\to 3$ BSM process we consider here gives rise to photons that are more energetic than those from the SM $\text{NC}1\gamma$ process, which arise from $Z$-boson exchange with the nucleus that induces a ground-state-to-ground-state transition in the nucleus with an associated photon PhysRevC.89.015503 . Therefore, this cross section constraint only weakly applies, with a $\sim$10% overlap in the true photon energy and angular distribution template in Ref. MicroBooNE:2025rsd to the coherent part of the $2\to 3$ process. We find that 10% of the coherent $2\to 3$ cross section satisfies the upper bound of $1.49\times 10^{-41}\leavevmode\nobreak\ \text{cm}^{2}$.
Existing searches from experiments such as NA62, PIENU, BaBar, CHARM, and E787 would constrain the mass and coupling of the mediators in our model. Since the $2\to 3$ scattering diagram involves an $\mathcal{O}(1)$ coupling between $Z^{\prime}$ and dark matter, the coupling $g_{Z^{\prime}}$ is constrained by invisible searches of $Z^{\prime}$ where $g_{Z^{\prime}}$ is the coupling of $Z^{\prime}$ with the SM sector. These searches include electron recoils at LSND, MiniBooNE, nuclear recoils at COHERENT COHERENT:2021pvd , CCM CCM:2021leg , missing energy at NA64 Banerjee:2019pds , and monophoton signals at BaBar BaBar:2014zli . Furthermore, the three-body decay constraints on charged pions and kaons at PIENU PIENU:2021clt ($\pi^{\pm}\to l^{\pm}\overline{\nu}_{l}Z^{\prime}$) and NA62 NA62:2021bji ($K^{\pm}\to\mu^{\pm}\overline{\nu}_{\mu}Z^{\prime}$) would also constrain the maximum allowed value of $g_{Z^{\prime}}$. Similarly, based on the exact Yukawa couplings between the scalar and SM fermions, $y_{\phi}$ is constrained from E787/E949 E787:2002qfb ; E787:2004ovg ; E949:2007xyy on the visible branching ratio of $K^{+}\to\pi^{+}+\phi$, MAMI A2atMAMI:2014zdf on the decay width of $\eta\to\pi^{0}+(\phi\to\gamma+\gamma)$, as well as PIENU ($\pi^{\pm}\to l^{\pm}\overline{\nu}_{l}\phi$) and NA62 ($K^{\pm}\to\mu^{\pm}\overline{\nu}_{\mu}\phi$).
The relevance of these bounds depends on the specifics of the UV model. For example, if the scalar does not couple directly to dark matter, but only via the $Z^{\prime}$, constraints from monophoton searches at BaBar and CHARM become relevant, which depend on the mass gap $m_{\phi}-2m_{\chi}$ in the decay $\phi\to\gamma\bar{\chi}\chi$, via an intermediate $Z^{\prime}$. Certain regions of the parameter space may be subject to supernova constraints (e.g., Dev:2020eam ), depending on additional model parameters that are not required to account for the anomaly. Additionally, our solutions require higher couplings and $m_{\chi}>10$ MeV; the solutions can bypass limits from supernova. The stellar cooling bounds do not constrain $m_{Z^{\prime}}$ above 1 MeV (e.g., Refs. Hardy:2016kme ; Li:2023vpv ). However, astrophysical models have more diverse and larger uncertainties Li:2023vpv . Since our solutions call for mediators and dark matter with masses greater than 10 MeV, constraints from CMB and BBN Giovanetti:2021izc are irrelevant.
Example UV complete models. We now discuss UV-complete $Z^{\prime}$ models which can lead to the fits in the Table 1. Since the process in Fig. 1 requires a new $Z^{\prime}$ and scalar, we consider a $U(1)_{T3R}$ gauge boson Dutta:2019fxn ; Dutta:2020scq that couples to second-generation quarks and third-generation leptons, and a scalar that couples to up, down, strange, and charm quarks. Therefore, the $s$ and $c$ quarks contribute to the dimension-5 coupling $g_{\phi Z^{\prime}\gamma}$, such that
$g_{\phi Z^{\prime}\gamma}y_{\phi}=\bigg{(}\frac{g_{T3R}y_{\phi}^{2}}{400}\bigg{)}\sum_{c,s}^{P=\{L,R\}}\frac{N_{c}Q_{q}x^{\phi}_{q_{P}}x^{T3R}_{q_{P}}}{m_{q}},$ (4)
where $g_{Z^{\prime}}$ is identified as $g_{T3R}$. The fractional gauge and scalar coupling fractions carried by fermions under $U(1)_{T3R}$ $Z^{\prime}$ and $\phi$ are denoted by $x_{q_{P}}^{T3R}$ and $x_{q_{P}}^{\phi}$, respectively. In the $U(1)_{T3R}$ case, the fractional gauge charges are $x^{T3R}_{s_{R}}=2,\leavevmode\nobreak\ x^{T3R}_{c_{R}}=-2,\leavevmode\nobreak\ x^{T3R}_{\tau_{R}}=2,$ and $x^{T3R}_{N_{R}}=-2$. This choice results in an anomaly-free $Z^{\prime}$ model. Here, the $T_{3R}$ quantum numbers do not contribute to the electric charge. The relevant fractional Yukawa couplings are $x^{\phi}_{s}=-1.016,\leavevmode\nobreak\ x^{\phi}_{c}=-0.3,\leavevmode\nobreak\ x^{\phi}_{u}=1,$ and $x^{\phi}_{d}=-1.2$. This choice strongly suppresses the branching fraction of $\phi\to\gamma\gamma$, satisfying the constraints from MAMI A2atMAMI:2014zdf .
Figure 3: Existing bounds on an invisible $U(1)_{T3R}$ gauge boson $Z^{\prime}$—i.e., $m_{Z^{\prime}}>2m_{\chi}$—that couples to second-generation quarks and third generation-leptons (top), and a scalar coupling to $u,d,s$, and $c$ quarks (bottom).
Figure 3 displays the existing bounds for the gauge boson and scalars for this model, assuming both to decay invisibly. We notice that for $m_{Z^{\prime}},m_{\phi}>80\leavevmode\nobreak\ \text{MeV}$, the most stringent bounds are set by NA62 that constrains the three-body decay of the charged kaon. Since constraints from $J/\Psi\to\gamma Z^{\prime}$ decays ParticleDataGroup:2024cfk are more suppressed than NA62 and those from $\Upsilon\to\gamma Z^{\prime}$ appear only at the loop-level, we refrain from including them in our analysis. Therefore, by adopting the maximum couplings allowed by the NA62 bounds to satisfy the dark matter fits in Table 1, we simultaneously maximize dark matter production from three-body decays into either $\phi$ or $Z^{\prime}$. Including the $s$ and $c$ quarks in the loop, we find that the benchmark points 2, 3, and 4 satisfy the requirements in Table 1, for $y_{\phi}$ and $g_{T3R}$ marked in Fig. 3. Note that the couplings simultaneously satisfy the constraints on scattering parameters, i.e., $g_{\phi Z^{\prime}\gamma}y_{\phi}$, as well as maximize the branching ratio of $K^{+}\to\mu^{+}+\nu_{\mu}+(Z^{\prime}/\phi)$, thereby allowing dark matter production to occur via the decays of both the $Z^{\prime}$ and $\phi$. These scalars can be produced through neutral mesons, via $\pi^{0},\eta\to\gamma\phi$, via operators such as $\partial_{\mu}\pi^{0}\partial_{\nu}\phi\tilde{F}^{\mu\nu}$ or $\partial_{\mu}\eta\partial_{\nu}\phi\tilde{F}^{\mu\nu}$. However, $\partial_{\mu}\pi^{0}\partial_{\nu}\phi\tilde{F}^{\mu\nu}$ and $\partial_{\mu}\eta\partial_{\nu}\phi\tilde{F}^{\mu\nu}$ that are suppressed by a cut-off scale square, i.e., $\frac{1}{\Lambda^{2}}$. Therefore, constraints from $\pi^{0}/\eta\to\gamma\phi$ are subdominant.
We note that in the scenario where $\chi$ is predominantly produced from $Z^{\prime}$, it need not couple to the scalar for our solution, thereby closing the scalar’s invisible decay modes. We discuss the bounds for visibly decaying scalars (i.e., $y_{\chi}=0$) in Appendix G.
@@
Conclusions. In this work, we demonstrated that light dark matter and/or neutrino scattering with a nucleus via light mediator exchange can account for the MiniBooNE excess in the $1\gamma 0p$ final state without invoking any additional neutrino species. These scenarios can be further probed in ongoing experiments such as ICARUS, SBND, and Belle-II. Importantly, the $2\to 3$ scattering mechanism plays a crucial role in explaining the excess: vector exchange with the nucleus tends to produce forward-peaked photons, while pseudoscalar exchange leads to off-forward photon emission due to the Lorentz structure of the interaction with the nucleus.
@@
Acknowledgements. We would like to thank Kevin J. Kelly, Rabi Mohapatra, Vishvas Pandey, and Austin Schneider for their helpful discussions. The work of BD and AK is supported by the U.S. DOE Grant DE-SC0010813. AT acknowledges support in part by the DOE grant DE-SC0010143. The work of RGV is supported by the Department of Energy Office of Science, High Energy Physics.